# Lifting the Veil on the Large Language Model Supply Chain: Composition, Risks, and Mitigations

### v2

@@ Abstract.
Large language models (LLM) have sparked significant impact with regard to both intelligence and productivity. In recent years, a great surge has been witnessed in the introduction of both commercial and open-source LLMs. Many businesses have adopted the LLMs into their applications to solve their own domain-specific tasks. However, integrating LLMs into specific business scenarios requires more than just utilizing the models themselves. Instead, it is a systematic process that involves substantial components, which are collectively referred to as the LLM supply chain. The LLM supply chain inherently carries risks. Therefore, it is essential to understand the types of components that may be introduced into the supply chain and the associated risks, enabling different stakeholders to implement effective mitigation measures. While some literature discusses risks associated with LLMs, there is currently no paper that clearly outlines the LLM supply chain from the perspective of both providing and consuming its components. As LLMs have become essential infrastructure in the new era, we believe that a thorough review of the LLM supply chain, along with its inherent risks and mitigation strategies, would be valuable for industry practitioners to avoid potential damages and losses, and enlightening for academic researchers to rethink existing approaches and explore new avenues of research. Our paper provides a comprehensive overview of the LLM supply chain, detailing the stakeholders, composing artifacts, and the supplying types. We developed taxonomies of risk types, risky actions, and mitigations related to various supply chain stakeholders and components. In summary, our work explores the technical and operational aspects of the LLM supply chain, offering valuable insights for researchers and engineers in the evolving LLM landscape.
@@ 1. Introduction
The large language models (LLMs) have sparked unprecedented discussion about the power of generative language models, showcasing their remarkable ability to understand human-written text, embody vast knowledge, and generate responses based on user instructions. In 2023 following ChatGPT’s debut, there has been a significant surge in the development and introduction of large language models, regarding both commercial LLMs and free open-source LLMs. Researchers and practitioners have extended the capabilities of LLMs beyond natural language processing, applying them in fields like software engineering, finance, and education, dramatically reshaping these areas (Science, 2024; Fan et al., 2023; bloomberg, 2024).
@@
Integrating LLMs involves more than just utilizing the models themselves. On one hand, although some applications may directly employ commercial LLMs, they still require extra solutions for pre-processing (e.g., Promptify (GitHub, 2024c) for prompt engineering, and GPTCache (GitHub, 2024b) for reducing LLM API call expenses), post-processing (e.g., filtering out unwanted content), integration (e.g., integrating with websites, servers, or mobile apps), and plugin interactions (e.g., Zapier (GitHub, 2024d) for linking with external apps). On the other hand, businesses may prefer open-source LLMs to enable flexible customizations, which demand more sophisticated components, during stages such as pre-training, model reusing, fine-tuning (e.g., Labelbox (labelbox, 2024)), model conversion (e.g., Onnx (onnx, 2024)), quantization (e.g., TensorRT (nvidia GitHub, 2024)) and deployment. Behind the directly used components, there exists a multitude of sub-components that are transitively depended upon, further increasing the overall dependency complexity. Furthermore, the supplying types in the LLM supply chain are more varied than in traditional software supply chains. Data and models introduce additional supply types, including partial data reuse, model parameter reuse through data augmentation, fine-tuning, etc. Overall, leveraging LLMs is a systematic process that entails numerous components and supplying types across multiple stages, collectively referred to as the LLM supply chain.
@@
Accordingly, the risks in the LLM supply chain have been significantly amplified due to an expanded attack surface and a growing number of participants. For stakeholders, it is essential to understand their providers and clients, how their artifacts are served, the risks associated with them, and the mitigation measures that are required. While the LLM supply chain shares similarities with the well-recognized open-source software supply chain (Ladisa et al., 2023), it also exhibits its uniqueness. For instance, attackers can exploit publicly accessible LLMs through prompt engineering to extract sensitive information (e.g., privacy leakage) or fine-tune open-source LLMs (eugeneyan, 2024) to embed malicious instructions within the model (e.g., backdoor attacks). Therefore, a comprehensive understanding of the types of components and supplying types in the LLM supply chain, thus to understand their corresponding risks is crucial to enable stakeholders to implement effective mitigation strategies.
While some literature discusses risks associated with LLMs, there is currently no paper that clearly outlines the LLM supply chain from the perspective of both providing and consuming its components. As a result, the inherent risks and potential mitigation strategies are less organized and comprehensive. Ladisa et al. (Ladisa et al., 2023) developed taxonomies for attack and defense techniques in software supply chains. However, their taxonomies do not account for LLM-specific components such as pre-trained LLMs, prompts, and data, making their survey less attractive for LLM supply chain stakeholders. Several surveys address the security and privacy challenges (Yao et al., 2024a; Cui et al., 2024b; Kumar et al., 2024; Yao et al., 2024a; Neel and Chang, 2023; Esmradi et al., 2023) and discuss the defense and mitigation strategies of LLMs. However, they did not explore which stakeholders and artifacts are affected, which may limit the ability to provide actionable insights and a clear scope for the relevant stakeholders. Additionally, Wang et al. (Wang et al., 2024b) were among the first to outline a research agenda for the LLM supply chain. They highlighted the key components in the LLM supply chain, illustrating them through the interactions within MLOps and DevOps. As the topic is still emerging and rapidly evolving, existing works may not comprehensively incorporate insightful ideas and perspectives from online preprints or web blogs. Therefore, we believe that a comprehensive LLM supply chain overview as well as a taxonomy of risks and clear guidance of mitigation in the LLM supply chain could be beneficial for both industrial and academic audiences. It could guide practitioners to pay attention to the LLM supply chain risks and provide a visionary map for academic researchers to devise new mitigation or defensive techniques.
@@
In our paper, we provide a comprehensive overview of the LLM supply chain, detailing the stakeholders, composing artifacts, and the supplying types, as illustrated in Figure 1. Our definition of the LLM supply chain differs from Want et al.’s work (Wang et al., 2024b) in emphasizing the LLM supply chain more on the basic components and their supply relationships between them. We developed taxonomies of risk types, risky actions, and mitigations related to various LLM supply chain stakeholders and components. Specifically, we fist holistically collected both academic and online resources. Then, we summarize the risk scenarios as stakeholders performing risky actions on specific supply chain components, which in turn lead to distinct risk types. We illustrate the risk scenarios from the perspective of upstream contributors, downstream users, and administrators. Furthermore, we list the types of various mitigation measures in response to the LLM supply chain risks.
@@
Our paper explores the technical and operational aspects of the LLM supply chain, offering a holistic understanding that draws valuable insights for researchers and engineers in software engineering, system architecture, software security, and data governance. The main contributions of our paper are as follows:
@@
  • We presented a comprehensive overview of the LLM supply chain, including artifacts, stakeholders, and supplying types.
@@
2. Large Language Model Supply Chain
We first provide our definition for the large language model supply chain. Then, we detail the stakeholders and composing components involved in the LLM supply chain, and provide a description of the LLM development stages. Lastly, we compare our work with existing surveys relevant to the LLM supply chain.
2.1. Definition
We present an overview of the large language model supply chain in Figure 1. While it may not capture the entire landscape exhaustively, we have aimed to make it as comprehensive as possible. The large language model (LLM) supply chain encompasses the ecosystem and sequence of processes integral to developing, training, deploying, and distributing applications that leverage large language models. This supply chain includes:
  • Stakeholders in various roles. Stakeholders take different roles based on the specific supply artifact and action. For instance, they may include developers contributing to the artifacts, organizations managing the platforms, or vendors providing cloud computing services.
  • Upstream artifacts in multiple forms. Upstream artifacts are those on which downstream applications have dependencies on. For example, these artifacts may include plaintext source code, binary executables, or serialized models.
  • Diverse supply relationships. As artifacts vary, the types of supply relationships differ. For instance, they may include augmentation relationships for data, cloning relationships for code, and merging relationships for models.
  • Development stages for various purposes. The development stages include a comprehensive scope for LLM application. e.g., data preprocessing, model pre-training, fine-tuning, integration, delivery, monitoring, and feedback.
2.2. Composing Components
We summarize five components in the LLM supply chain, i.e., stakeholder, artifact, toolchain, platform, and LLM applications (apps).
Figure 1. Overview of the Large Language Model Supply Chain
2.2.1. Stakeholders
We summarize four roles of stakeholders in the LLM supply chain, i.e., contributors, consumers, administrators, and user. Note that we categorize the stakeholders as roles because an individual can occupy a single role or assume multiple roles simultaneously, such as being both an contributor and an administrator in an open-source package platform.
Contributors. We collectively refer to contributors engaged in development activities and sharing artifacts, such as data, models, prompts, or third-party libraries, as contributors. Their roles vary depending on the types of artifacts they contribute.
  • Platform Contributors are responsible for contributing to hosting platforms (e.g., hosting artifacts), hardware platforms, and serving platforms (i.e., Google Cloud (Cloud, 2024), Amazon Web Services (Services, 2024)). They are obligated to provided secure, stable, isolated, and managed environments. i.e., malicious content detection, safe system images, authorized cloud accounts, etc.
  • Artifact Contributors collect, process, and upload various forms of data, code, prompts, and models to the hosting platforms. They can be data engineers, model developers, prompt engineers, and library developers.
  • Toolchain Contributors are responsible for developing, maintaining, and improving the tools used throughout the model lifecycle. e.g., Jupyter, TensorFlow for model development, Jenkins, GitHub Actions in CI/CD.They ensure that these tools are reliable, scalable, and compatible with the requirements from Toolchain Consumers.
Consumers. We distinguish consumers depending on their consuming components, including platform consumers, artifact consumers, toolchain consumers, and app consumers.
  • Platform Consumers are entities that use the platform infrastructure for various purposes, such as accessing hosting services, utilizing computational resources, or interacting with serving platforms for deploying and running models and applications.
  • Artifact Consumers are those who utilize the artifacts produced and shared within the platform. This includes data sets, code, prompts, and models.
  • Toolchain Consumers are individuals or teams that leverage the toolchain components for developing, deploying, and maintaining models. They use model development, CI/CD, and deployment tools to streamline the lifecycle of models and applications.
  • App Consumers are end-users or other systems that interact with applications built using the platform’s resources and artifacts.
Administrators. We summarize administrators as platform administrators and apps administrators.
  • Platform Administrator include two sub roles. The platform administrators manage and oversee the infrastructure, ensuring the system’s operational integrity, security, and performance. The project administrators, on the other hand, focus on managing repositories of artifacts, such as data repositories, model repositories, prompt repositories/websites, and library repositories. They hold the privileges to merge pull requests, manage versioning, and onboard new contributors.
  • Apps Administrator is responsible for managing the lifecycle of applications that rely on the artifacts and models within the platform. This includes deploying, configuring, and monitoring applications.
2.2.2. Artifacts
The LLM supply chain artifacts include data, model, prompts and code.
Data. Data serves as the foundational element in the workflow. It can be augmented and preprocessed from the original data to improve its quality, ensuring it is ready for use in training and fine-tuning the model. Retrieval methods bring in relevant data from various sources, facilitating the prompt generation process. They can be used as the data for pre-training or fine-tuning, or as the knowledge database for retrieval augmented generation.
Model. The model represents the large language model that is trained and fine-tuned using the processed data and instructions from the prompt. The model may be adjusted or modified by splitting, merging, serializing, or deserializing components as needed to optimize performance. It generates predictions or outputs based on the prompts used as the instructs.
Prompt. The prompt is crafted to instruct the model effectively, guiding it on what information to retrieve, process, or generate. Prompts can be refined to adapt to specific scenarios, leading to more accurate or tailored instructions that improve model outputs.
Code. The code encompasses the software and scripts utilized for data processing, model training, and prompt refinement. This includes code that directly invokes relevant library APIs, as well as the libraries themselves as dependencies. The code can be cloned for customization to enhance compatibility and functionality.
2.2.3. Platforms
Hosting Platform. These platforms facilitate the storage, sharing, and versioning of software packages, models, data, and related resources required for LLM development and deployment. e.g., Data Hosting Platforms (e.g., Kaggle (LLC, 2024) for datasets), Model Hosting Platforms (e.g., Hugging Face (Face, 2024) for pre-trained models), Repository Hosting Platforms (e.g., GitHub (GitHub, 2024a) for code repositories), Package Registries (e.g., PyPI (Foundation, 2024) for Python packages, Debian (Project, 2024) for Linux packages), Plugin Marketplaces (e.g., Visual Studio Marketplace (Microsoft, 2024) for development tools), and Prompt Sharing Platforms where users can share and discover prompt engineering techniques.
Hardware Platform. These platforms consist of specialized hardware designed to optimize the performance of LLMs. They include dedicated chips and accelerators (e.g., NVIDIA GPUs, Google TPUs) that provide the computational power necessary to handle large datasets and complex models, reducing training time and improving efficiency in deployment.
Serving Platform. These platforms offer cloud-based services to support the training, deployment, and maintenance of LLMs. They include Cloud Computing Platforms that provide scalable computing resources, such as Google Cloud (Cloud, 2024) and Amazon Web Services (Services, 2024). These platforms allow users to leverage high-performance infrastructure, including GPUs and TPUs, for intensive tasks like model training and inference.
2.2.4. Toolchain
We discuss the relevant toolchain with regard to model development toolchain, model deployment toolchain, and CI/CD toolchain.
Model Development Toolchain. Tools used for training or fine-tuning the LLMs. If an attacker implants malicious code in these tools, it could lead to corrupted models or abnormal outputs. Environment Image is the base environment (e.g., Docker containers) in which the model runs could be compromised, affecting system security or performance. Model Serializer and Deserializer save or load models. If compromised, they could modify the model’s internal state or introduce backdoors.
Model Deployment Toolchain. Tools specifically used for creating and deploying LLMs. Any manipulation or vulnerabilities here could compromise the model’s behavior or safety. Model Quantization Tool reduce the size of the deployment model, making it more efficient. If compromised, these tools could degrade model performance or introduce malicious behavior. Model Conversion Tool converts models between formats or versions. A poisoned conversion process could lead to abnormal model behavior or failures.
CI/CD Toolchain. Continuous Integration/Continuous Deployment systems that automate the software delivery process. A compromised CI/CD pipeline could introduce vulnerabilities during model updates.
2.2.5. Applications
2.3. Model Development Lifecycle
Figure 2. Large Language Model Development Stages
Figure 2 illustrates the four essential stages of the large language model (LLM) lifecycle. Each stage includes specific processes crucial to building, deploying, and maintaining LLMs effectively. In the Preparation stage, foundational work is completed, including data pre-processing to clean and organize raw data, foundation model (FM) preparation to set up initial models, and prompt engineering to guide model outputs effectively. These steps ensure that the inputs are optimized for subsequent stages. The Training stage focuses on developing the model’s capabilities. Here, model pre-training introduces the model to large datasets, enabling it to understand language patterns. This is followed by fine-tuning on specific tasks to enhance accuracy and model compression to optimize the model for efficient deployment. In the Deployment stage, the trained model is integrated into real-world applications. This involves model integration to embed the LLM within applications, model conversion to make it compatible with different systems, and model serving to set up scalable infrastructure for user interactions. Finally, the Delivery stage ensures the ongoing performance and governance of the model. Model monitoring is conducted to track performance and detect issues, model maintenance includes regular updates and retraining to retain effectiveness, and model governance oversees the ethical and regulatory aspects, ensuring compliance and responsible usage.
2.4. Related Surveys
@@
We conducted a survey of the security aspects of large language models (LLMs). Specifically, we gathered 12 surveys from our literature collection (see Section 3.1) that discuss various security-related topics, including LLM risks, privacy challenges, adversarial attacks, and defenses. For ease of comparison, we categorized the topics into two main areas: attacks and risks and defensive techniques. Additionally, we unified the terminology and definitions across the surveys to enable a clearer comparison.
@@
Attack and Risks. In the context of attacks and risks, the Output Risk is extensively covered in eight survey papers, which refers to the potential for LLMs to generate harmful, untruthful, hallucinatory, or unhelpful content. Following this, Privacy Attacks are another major concern, with nine papers addressing various aspects such as membership inference attacks, privacy leakage (e.g., PPI) attacks, gradient leakage attacks, model inversion attacks, and attribute inference attacks. Besides, Prompts Attack is covered in 7 papers, including prompt injection attack and jailbreaking attack. Unfortunately, only a few papers (Pankajakshan et al., 2024; Cui et al., 2024b) mention Toolchain Attacks. Yao et al. (Yao et al., 2024a) discuss Supply Chain Vulnerabilities, and Wu et al. (Wu et al., 2024) highlight issues like malicious webtool misuse and cross-session access, which are closely related but account for a subset of Toolchain Attacks. Therefore, the topic of toolchain attacks was not explored in depth. Finally, only Neel et al. (Neel and Chang, 2023) addressed the issue of Copyright Risks.
@@
Defenses and Mitigation. Defenses and Mitigation are suggested in accordance with the corresponding attacks and risks. For Prompt Attacks, existing works propose Input Sanitization (Pankajakshan et al., 2024; Iqbal et al., 2023). To mitigate Output Risk, approaches like Output Sanitization are recommended by several studies (Yao et al., 2024a; Cui et al., 2024b; Kumar et al., 2024). For Data Attacks, the literature suggests techniques such as Data Cleaning (Yao et al., 2024a), Data De-duplication (Neel and Chang, 2023), and Data Sanitization (Esmradi et al., 2023). However, these defensive techniques are not well-summarized across the surveyed works. While some methods, such as Differential Privacy, Federated Learning, or LLM Alignment, refer to specific technical measures designed to defend against particular attacks or mitigate associated risks, other defensive measures are less detailed. For instance, strategies like Output Detection and Input Sanitization are suggested to address Output Risk and Prompt Attacks, respectively, but they lack in-depth technical details on how these defenses are implemented.
@@
Comparison to Existing Surveys. Our paper contributes the following contributions which are novel and incremental to existing surveys. First, we conducted a comprehensive survey in both academic literature and online resources regarding the open-source LLM supply chain. Second, we provide a comprehensive overview and highlight the key stakeholders and components in the LLM supply chain. Third, we summarize the risk scenarios as stakeholders performing risk actions on specific supply chain components, as well as risk types. We illustrate the risk scenarios from the perspective of three major stakeholders. Furthermore, we list the types of various mitigation measures in response to the LLM supply chain risks.

### v3

# Lifting the Veil on Composition, Risks, and Mitigations of the Large Language Model Supply Chain

@@ Abstract.
Large language models (LLMs) have sparked significant impact with regard to both intelligence and productivity. Numerous enterprises have integrated LLMs into their applications to solve their own domain-specific tasks. However, integrating LLMs into specific scenarios is a systematic process that involves substantial components, which are collectively referred to as the LLM supply chain. A comprehensive understanding of LLM supply chain composition, as well as the relationships among its components, is crucial for enabling effective mitigation measures for different related risks. While existing literature has explored various risks associated with LLMs, there remains a notable gap in systematically characterizing the LLM supply chain from the dual perspectives of contributors and consumers. In this work, we develop a structured taxonomy encompassing risk types, risky actions, and corresponding mitigations across different stakeholders and components of the supply chain. We believe that a thorough review of the LLM supply chain composition, along with its inherent risks and mitigation measures, would be valuable for industry practitioners to avoid potential damages and losses, and enlightening for academic researchers to rethink existing approaches and explore new avenues of research.
@@ 1. Introduction
The large language models (LLMs) have sparked unprecedented discussion about the power of generative language models. There has been a significant surge in the development and introduction of large language models, regarding both commercial LLMs and free open-source LLMs. Researchers and practitioners have extended the capabilities of LLMs beyond natural language processing, applying them in fields like software engineering, finance, and education, dramatically reshaping these areas (Science, 2024; Fan et al., 2023; bloomberg, 2024).
@@
Integrating LLMs involves more than just utilizing the models themselves. On the one hand, although some applications may directly employ commercial LLMs, they still require additional solutions such as pre-processing (e.g., Promptify (GitHub, 2024c) for prompt engineering, GPTCache (GitHub, 2024b) for reducing LLM API call expenses), etc. On the other hand, businesses may prefer open-source LLMs to enable flexible customizations, which demand more sophisticated components, during stages such as fine-tuning (e.g., Labelbox (labelbox, 2024)), model conversion (e.g., Onnx (onnx, 2024)), quantization (e.g., TensorRT (nvidia GitHub, 2024)), etc. Consequently, there exists a multitude of sub-components that are transitively depended upon. Furthermore, the composition types in the LLM supply chain are more varied than in the traditional software supply chain. Overall, leveraging LLMs is a systematic process that entails numerous components and composition types, collectively referred to as the LLM supply chain.
@@
Accordingly, we are confronted with a fresh LLM supply chain ecosystem, where the attack surface has expanded but remains understudied, and the number of participants is rapidly increasing. While the LLM supply chain shares similarities with the well-recognized open-source software supply chain (Ladisa et al., 2023), it also exhibits its uniqueness. For instance, attackers can exploit publicly accessible LLMs to extract sensitive information (e.g., privacy leakage) or fine-tune open-source LLMs (eugeneyan, 2024) to embed malicious instructions (e.g., backdoor attacks).
Figure 1. Overview of Large Language Model Supply Chain
We provide our definition for the large language model supply chain, as shown in Fig. 1. It encompasses the ecosystem and sequence of processes integral to developing, training, deploying, and distributing applications that leverage large language models, including:
  • Stakeholders in various roles. e.g., developers contributing to the artifacts, organizations managing the platforms, or vendors providing cloud computing services.
  • Upstream artifacts in multiple forms. e.g., plaintext source code, binary executables, or serialized models.
  • Diverse supply relationships. e.g., augmentation relationships for data, cloning relationships for code, and merging relationships for models.
  • Development stages for various purposes. e.g., data preprocessing, model pre-training, fine-tuning, integration, and monitoring.
While existing literature discusses risks associated with LLMs, there rarely exists a comprehensive work that outlines the LLM supply chain from the perspective of contributors and consumers. We aim to answers the following research questions:
RQ1 Composition Analysis: What new components accompanied with their defining characteristics, emerge in the LLM supply chain, compared with the software supply chain?
RQ2 Risk Analysis: How to characterize the risks posed by the LLM supply chain? What are their uniqueness, compared with the software supply chain?
RQ3 Mitigation Analysis: How to mitigate the risks in the LLM supply chain?
@@
To study RQ1, we provide a comprehensive overview of the LLM supply chain, detailing the stakeholders, composing artifacts, and the supplying types, as illustrated in Figure 1. Our definition of the LLM supply chain differs from Want et al.’s work (Wang et al., 2024b) in emphasizing the LLM supply chain more on the basic components and their supply relationships between them. To explore RQ2, we developed taxonomies of risk types, risky actions, and mitigations related to various LLM supply chain stakeholders and components. Specifically, we fist holistically collected both academic and online resources. Then, we summarize the risk scenarios as stakeholders performing risky actions on specific supply chain components, which in turn lead to distinct risk types. We illustrate the risk scenarios from the perspective of upstream contributors, downstream users, and administrators. Furthermore, to study RQ3, we list the types of various mitigation measures in response to the LLM supply chain risks.
@@
Our paper explores the technical and operational aspects of the LLM supply chain, offering a holistic understanding that draws valuable insights for researchers and engineers in software engineering, system architecture, software security, and data governance. The contributions of our paper are as follows:
@@
  • We presented a comprehensive overview of the LLM supply chain, including artifacts, stakeholders, and composition types.
@@
2. Related Work
@@
Various security-related aspects of large language models are discussed in existing work. For ease of comparison, we focus our topics in attacks and risks and defensive techniques. Additionally, we unified the terminology and definitions across the surveys to enable a clearer comparison.
@@
Attack and Risks. In the context of attacks and risks, the Output Risk is extensively covered in eight survey papers, which refers to the potential for LLMs to generate harmful, untruthful, hallucinatory, or unhelpful content. Following this, Privacy Attacks are another major concern, with nine papers addressing various aspects such as membership inference attacks, privacy leakage (e.g., PPI) attacks, gradient leakage attacks, model inversion attacks, and attribute inference attacks. Besides, Prompts Attack is covered in seven papers, including prompt injection attack and jailbreaking attack. Unfortunately, only a few papers (Pankajakshan et al., 2024; Cui et al., 2024b) mention Toolchain Attacks. Yao et al. (Yao et al., 2024) discuss Supply Chain Vulnerabilities, and Wu et al. (Wu et al., 2024) highlight issues like malicious webtool misuse and cross-session access, which are closely related but account for a subset of Toolchain Attacks. Therefore, the topic of toolchain attacks was not explored in depth. Finally, only Neel et al. (Neel and Chang, 2023) addressed the issue of Copyright Risks.
@@
Defenses and Mitigation. Defenses and Mitigation are suggested in accordance with the corresponding attacks and risks. For Prompt Attacks, existing works propose Input Sanitization (Pankajakshan et al., 2024; Iqbal et al., 2023). To mitigate Output Risk, approaches like Output Sanitization are recommended by several studies (Yao et al., 2024; Cui et al., 2024b; Kumar et al., 2024). For Data Attacks, the literature suggests techniques such as Data Cleaning (Yao et al., 2024), Data De-duplication (Neel and Chang, 2023), and Data Sanitization (Esmradi et al., 2023). However, these defensive techniques are not well-summarized across the surveyed works. While some methods, such as Differential Privacy, Federated Learning, or LLM Alignment, refer to specific technical measures designed to defend against particular attacks or mitigate associated risks, other defensive measures are less detailed. For instance, strategies like Output Detection and Input Sanitization are suggested to address Output Risk and Prompt Attacks, respectively, but they lack in-depth technical details on how these defenses are implemented.
@@
Comparison to Existing Surveys. Our paper contributes the following contributions which are novel and incremental to existing surveys. First, we conducted a comprehensive survey in both academic literature and online resources regarding the open-source LLM supply chain. Second, we provide a comprehensive overview and highlight the key stakeholders and components in the LLM supply chain. Our definition of the LLM supply chain differs from Want et al.’s work (Wang et al., 2024b) in emphasizing the LLM supply chain more on the basic components and their supply relationships between them. Third, we summarize the risk scenarios as stakeholders performing risky actions on specific supply chain components, as well as risk types. We illustrate the risk scenarios from the perspective of three major stakeholders. Furthermore, we list the types of various mitigation measures in response to the LLM supply chain risks.